In a nuclear reactor fuel consumption is 1 ug/sec then power generated is?? 

Grade:10

1 Answers

Arun
5 years ago
Dear Ankit
 
Power generated  p = mc^2 / t
 
p = 1 * 10^-6 * 9 * 10^16
 
p = 9 * 10^7 K watt
 
Regards
Arun (askIITians forum expert)